OpenStudy (anonymous):

what is the length of the missing side if the other two sides are 3 and 5? please help i gotta get 15 6 assignments done by monday!

Parth (parthkohli):

I believe you have a diagram, or the number of sides in the figure.

OpenStudy (anonymous):

what ?

Parth (parthkohli):

It'd be even better if it has a specific name like 'square', 'rectangle'.

Parth (parthkohli):

Is there a diagram in this question?

OpenStudy (lgbasallote):

use the pythagorean theorem \[a = \sqrt{c^2 - b^2}\] \[a = \sqrt{5^2 - 3^2}\] i think this your problem right?

Parth (parthkohli):

Hmm. If it is a right triangle, then we don't know if 5 is the hypotenuse or the longer leg.

OpenStudy (anonymous):

triangle

Parth (parthkohli):

But, I go for \(5 \text{ is the hypotenuse, because we have a triplet}\) with 5 as the hypotenuse.

Parth (parthkohli):

\( \color{Black}{\Rightarrow 3^2 + b^2 = 5^2}\) \( \color{Black}{\Rightarrow b^2 = 5^2 - 3^2 }\) \( \color{Black}{\Rightarrow b = \sqrt{5^2 - 3^2}}\)

OpenStudy (anonymous):

5 is the base and 3 is the up and down side

Parth (parthkohli):

Can you draw it out for us?

OpenStudy (anonymous):